OpenStudy (anonymous):

Solve for variable. 2(5x-2)=2(15x+14). Can someone show me how to solve this.

OpenStudy (anonymous):

Is anyone able to help with this question?

OpenStudy (anonymous):

X=-1.6

OpenStudy (anonymous):

Can you show me how you got that

OpenStudy (anonymous):

The way this works is divide both sides by 2. You are left with: \[5x-2=15x+14\] Then gather like terms: \[-2-14=15x-5x\] Then simplify: \[-16=10x\] Then isolate for x by dividing both sides by 10 \[x=-1.6\]

OpenStudy (anonymous):

I'm still not understanding. Can someone explain it a little better. With more steps.